# Active Learning of Model Discrepancy with Bayesian Experimental Design

### v1

@@ Abstract
Digital twins have been actively explored in many engineering applications, such as manufacturing and autonomous systems. However, model discrepancy is ubiquitous in most digital twin models and has significant impacts on the performance of using those models. In recent years, data-driven modeling techniques have been demonstrated promising in characterizing the model discrepancy in existing models, while the training data for the learning of model discrepancy is often obtained in an empirical way and an active approach of gathering informative data can potentially benefit the learning of model discrepancy. On the other hand, Bayesian experimental design (BED) provides a systematic approach to gathering the most informative data, but its performance is often negatively impacted by the model discrepancy. In this work, we build on sequential BED and propose an efficient approach to iteratively learn the model discrepancy based on the data from the BED. The performance of the proposed method is validated by a classical numerical example governed by a convection-diffusion equation, for which full BED is still feasible. The proposed method is then further studied in the same numerical example with a high-dimensional model discrepancy, which serves as a demonstration for the scenarios where full BED is not practical anymore. An ensemble-based approximation of information gain is further utilized to assess the data informativeness and to enhance learning model discrepancy. The results show that the proposed method is efficient and robust to the active learning of high-dimensional model discrepancy, using data suggested by the sequential BED. We also demonstrate that the proposed method is compatible with both classical numerical solvers and modern auto-differentiable solvers.
@@ 1 Introduction
Data-driven modeling techniques have been demonstrated promising in many science and engineering applications [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. Considering that data informativeness serves as a critical component in all data-driven modeling techniques, an active approach to identifying the most informative data has the potential to further advance state-of-the-art data-driven modeling. Experimental design provides such a systematical approach for the active acquisition of data, which has been demonstrated effective in various types of problems, such as the estimation of some unknown model parameters [15, 16, 17, 18, 19]. However, classical experimental design via solving an optimization problem is limited by its propensity to fall into local optima when applied to nonlinear models and its inability to fully incorporate prior knowledge [20].
Bayesian experimental design (BED) addresses these limitations by utilizing a probabilistic framework that incorporates prior information and updates parameter distributions as new data are acquired [20, 21]. This approach not only enhances the robustness of the design and data acquisition process but also provides a more flexible and adaptive framework that can efficiently handle complex, nonlinear models [22, 23]. Foundational work [24, 25] introduced Bayesian decision theory into the field of experimental design, emphasizing the importance of selecting a utility function that appropriately reflects the purpose of the experiment, which is critical for identifying optimal designs [26], and various popular criteria used to find optimal designs were summarized in [27]. The growing popularity of sequential or adaptive designs in Bayesian design literature highlights their flexibility and efficiency over static designs [28, 29, 30]. Numerous novel strategies and techniques have been developed to overcome the computational challenges for Bayesian optimal design problems [31, 32, 33, 34], and a comprehensive review of Bayesian OED for its formulations and computations can be found in [35].
@@ 2 Methodology
The model discrepancy is ubiquitous in the modeling and simulation of complex dynamical systems and poses a key challenge to the performance of Bayesian optimal experimental design. Unlike the unknown parameters of an existing physics-based model that are often in a low-dimensional vector space, the model discrepancy is in infinite-dimensional function space, for which the approximation approaches (e.g., based on neural networks) can involve unknown coefficients in a high-dimensional vector space and thus make full Bayesian approach such as MCMC practically infeasible. The general goal of this work is to leverage the optimal data identified initially for interested physical parameters by the standard BED via an ensemble-based method and to efficiently learn the unknown coefficients of the neural network that approximates the model discrepancy.
@@
where the system state $\mathbf{u}(\mathbf{z},t)$ is often a spatiotemporal field for many engineering applications, and $\mathcal{G}^{\dagger}$ denotes the true dynamics. If the detailed form of true dynamics is known, the BED can be used straightforwardly to estimate the unknown parameters $\bm{\theta}^{\dagger}$. In practice, the form of true dynamics is often unknown for many complex dynamical systems. Assuming that a physics-based model $\mathcal{G}(\mathbf{u};\bm{\theta}_{\mathcal{G}})$ exists to approximate the true dynamics $\mathcal{G}^{\dagger}$, we focus on the learning of model discrepancy $\mathcal{G}^{\dagger}-\mathcal{G}$, which is characterized by a neural network $\textrm{NN}(\mathbf{u};\bm{\theta}_{\textrm{NN}})$ and leads to the modeled system:
@@
Using a neural-network-based model to characterize the model discrepancy has recently been explored with different choices of neural networks, e.g., neural operators [53] and diffusion models [54], with a more general review by [55]. In the context of BED, the design $\mathbf{d}$ corresponds to the spatiotemporal coordinate $(\mathbf{z},t)$. The measurement $\mathbf{y}$ is the observation of variable $\mathbf{u}$ that represents the state of the physical system at the design $\mathbf{d}$, i.e., $\mathbf{y}=\mathbf{u}(\mathbf{d})+\bm{\eta}$, where $\bm{\eta}\sim\mathcal{N}(\bm{0},\bm{\Gamma})$ represents the measurement noise which is assumed to be Gaussian in this work when observing the state from the true system.
@@ 2.1 Bayesian Experimental Design
where $\mathbf{y}\in\mathcal{Y}\in\mathbb{R}^{d_{\mathbf{y}}}$ is data from the design $\mathbf{d}\in\mathcal{D}$, $\bm{\theta}\in\bm{\Theta}$ denotes the target parameters, and $U$ is the utility function taking the inputs of $\mathbf{y}$ and $\bm{\theta}$ and returning a real value. The optimal design $\mathbf{d}^{*}$ is obtained by maximizing the expected utility function $\mathbb{E}[U(\bm{\theta},\mathbf{y},\mathbf{d})]$ over the design space $\mathcal{D}$. The term $p(\bm{\theta},\mathbf{y}|\mathbf{d})$ is the joint conditional distribution of data and parameters. $p(\mathbf{y}|\bm{\theta},\mathbf{d})$ is the likelihood. One thing that should be noted is that $p(\bm{\theta}|\mathbf{d})$ is the prior distribution of parameters, which is often assumed to be independent of the design $\mathbf{d}$: $p(\bm{\theta}|\mathbf{d})=p(\bm{\theta})$.
The utility function $U(\bm{\theta},\mathbf{y},\mathbf{d})$ can be regarded as the reward obtained from the data $\mathbf{y}$ for the corresponding design $\mathbf{d}$. This work employs Kullback-Leibler divergence between the posterior and the prior distribution of parameters $\bm{\theta}$:
@@ 2.2 Model Discrepancy Correction
More specifically, the iterative approach at $i$-th stage starts with solving a BED problem for $\bm{\theta}_{\mathcal{G}}$ based on the current estimation of $\bm{\theta}_{\text{NN}}$ to obtain the optimal design $\mathbf{d}$ in Eq. (7). With the fixed design $\mathbf{d}$ and the neural network parameters $\bm{\theta}_{\text{NN}}$, we further obtain the maximum a posterior (MAP) point $\bm{\theta}_{\mathcal{G}}^{*}$:
$\bm{\theta}_{\mathcal{G}}^{*}=\operatorname*{arg\,max}_{\bm{\theta}_{\mathcal{G}}}\{p(\mathbf{y}|\bm{\theta}_{\mathcal{G}},\mathbf{d};\bm{\theta}_{\text{NN}})p(\bm{\theta}_{\mathcal{G}};\bm{\theta}_{\text{NN}})\}.$ (8)
With the MAP estimation of $\bm{\theta}_{\mathcal{G}}^{*}$ further being fixed, the objective function $L(\bm{\theta}_{\text{NN}};\ \bm{\theta}_{\mathcal{G}}^{*},\mathbf{y},\mathbf{d})$ is defined by the likelihood function $p(\mathbf{y}|\bm{\theta}_{\mathcal{G}}^{*},\mathbf{d};\bm{\theta}_{\text{NN}})$ , and the updated estimation of $\bm{\theta}_{\text{NN}}$ can be obtained via:
@@
We omit the stage index $i$ in this section for simplicity in notation. In summary, there are three key steps at each stage of the hybrid framework illustrated in Fig. 1: (i) performing BED to obtain optimal design $\mathbf{d}$, (ii) obtaining the MAP estimation $\bm{\theta}_{\mathcal{G}}^{*}$, and (iii) updating the estimation of $\bm{\theta}_{\text{NN}}$. A sketch of proof for the existence of a globally optimal solution is presented in B.
In practice, we sample $m$ points for the estimation of $\bm{\theta}_{\mathcal{G}}^{*}$ to enhance the robustness of the proposed algorithm:
$\{\bm{\theta}_{\mathcal{G}}^{(1)},\bm{\theta}_{\mathcal{G}}^{(2)},...,\bm{\theta}_{\mathcal{G}}^{(m)}\}=\operatorname{arg\,top}_{m}p(\mathbf{y}|\bm{\theta}_{\mathcal{G}},\mathbf{d};\bm{\theta}_{\text{NN}})p(\bm{\theta}_{\mathcal{G}};\bm{\theta}_{\text{NN}}),$ (10)
where $\operatorname{arg\,top}_{m}$ denotes the process that identifies the set of parameters corresponding to the top $m$ largest values of a function, and the $\bm{\theta}_{\mathcal{G}}^{*}$ in Eq. (9) is approximated by those samples (more related to posterior mean) :
@@
It should be noted that $\mathrm{d}\mathbf{u}/\mathrm{d}\bm{\theta}_{\text{NN}}$ is often expensive to evaluate directly when $\mathbf{u}$ and $\bm{\theta}_{\text{NN}}$ are both high-dimensional. Therefore, efficient methods have been developed to obtain $\mathrm{d}\text{Obj}/\mathrm{d}\bm{\theta}_{\text{NN}}$ such as automatic differentiation [60, 61] or adjoint methods [62, 63].
As automatic differentiation is essentially a discretized version of the adjoint method, we only briefly introduce the adjoint method here. To avoid directly calculating $\mathrm{d}\mathbf{u}/\mathrm{d}\bm{\theta}_{\text{NN}}$, an adjoint state $\lambda(t)$ is defined such that $\lambda(t)^{\top}=\partial L/\partial\mathbf{u}(t)$. For the modeled system in Eq. (2), the adjoint state in time range $[t_{0},t_{1}]$ can be obtained by backwardly solving:
$\frac{\mathrm{d}\lambda(t)}{\mathrm{d}t}&=-\left(\frac{\partial(\mathcal{G}+\text{NN})}{\partial\mathbf{u}}\right)^{\top}\lambda(t)\\ \lambda^{\top}(t_{1})&=\frac{\partial L}{\partial\mathbf{u}(t_{1})}\end{aligned}\end{array}\right.,$ (13)
where NN$/\partial\mathbf{u}∂ N$ can be obtained directly via the backpropagation of the neural network NN and $\partial\mathcal{G}/\partial\mathbf{u}$ is the Jacobian of the physics-based model $\mathcal{G}$. In practice, the Jacobian can be derived analytically or approximated by the coefficient matrix from the numerical solver, which often formulates a linear system. With the adjoint state obtained, the gradient $\mathrm{d}L/\mathrm{d}\bm{\theta}_{\text{NN}}$ can then be calculated as follows:
[…]

### v2

@@ Abstract
Digital twins have been actively explored in many engineering applications, such as manufacturing and autonomous systems. However, model discrepancy is ubiquitous in most digital twin models and has significant impacts on the performance of using those models. In recent years, data-driven modeling techniques have been demonstrated promising in characterizing the model discrepancy in existing models, while the training data for the learning of model discrepancy is often obtained in an empirical way and an active approach of gathering informative data can potentially benefit the learning of model discrepancy. On the other hand, Bayesian experimental design (BED) provides a systematic approach to gathering the most informative data, but its performance is often negatively impacted by the model discrepancy. In this work, we build on sequential BED and propose an efficient approach to iteratively learn the model discrepancy based on the data from the BED. The performance of the proposed method is validated by a classical numerical example governed by a convection-diffusion equation, for which full BED is still feasible. The proposed method is then further studied in the same numerical example with a high-dimensional model discrepancy, which serves as a demonstration for the scenarios where full BED is not practical anymore. In this example with time-varying velocity, the network only learns the stationary discrepancy and leaves the time dependence being handled by the physics-based model itself, which helps mitigate ill-posedness of the joint learning problem. An ensemble-based approximation of information gain is further utilized to assess the data informativeness and to enhance learning model discrepancy. The results show that the proposed method is efficient and robust to the active learning of high-dimensional model discrepancy, using data suggested by the sequential BED. We also demonstrate that the proposed method is compatible with both classical numerical solvers and modern auto-differentiable solvers.
@@ 1 Introduction
Data-driven modeling techniques have been demonstrated promising in many science and engineering applications [1, 2, 3], with various types of methods being developed such as model reduction [4, 5], physics-informed machine learning [6, 7, 8], operator learning [9, 10, 11], causal learning [12, 13], and data assimilation [14, 15, 16]. Considering that data informativeness serves as a critical component in all data-driven modeling techniques, an active approach to identifying the most informative data has the potential to further advance state-of-the-art data-driven modeling. Experimental design provides such a systematical approach for the active acquisition of data, which has been demonstrated effective in various types of problems, such as the estimation of some unknown model parameters [17, 18, 19, 20]. However, classical experimental design via solving an optimization problem is limited by its propensity to fall into local optima when applied to nonlinear models and its inability to fully incorporate prior knowledge [21].
Bayesian experimental design (BED) addresses these limitations by utilizing a probabilistic framework that incorporates prior information and updates parameter distributions as new data are acquired [21, 22]. This approach not only enhances the robustness of the design and data acquisition process but also provides a more flexible and adaptive framework that can efficiently handle complex, nonlinear models [23, 24]. Foundational work [25, 26] introduced Bayesian decision theory into the field of experimental design, emphasizing the importance of selecting a utility function that appropriately reflects the purpose of the experiment, which is critical for identifying optimal designs [27], and various popular criteria used to find optimal designs were summarized in [28]. The growing popularity of sequential or adaptive designs in Bayesian design literature highlights their flexibility and efficiency over static designs [29, 30, 31]. Numerous novel strategies and techniques have been developed to overcome the computational challenges for Bayesian optimal design problems [32, 33, 34, 35], and a comprehensive review of BED for its formulations and computations can be found in [36].
@@ 2 Methodology
The model discrepancy is ubiquitous in the modeling and simulation of complex dynamical systems [54] and poses a key challenge to the performance of Bayesian optimal experimental design. Unlike the unknown parameters of an existing physics-based model that are often in a low-dimensional vector space, the model discrepancy is in infinite-dimensional function space, for which the approximation approaches (e.g., based on neural networks) can involve unknown coefficients in a high-dimensional vector space and thus make full Bayesian approach such as MCMC practically infeasible. The general goal of this work is to leverage the optimal data identified initially for interested physical parameters by the standard BED via an ensemble-based method and to efficiently learn the unknown coefficients of the neural network that approximates the model discrepancy.
@@
where the system state $\mathbf{u}(\mathbf{z},t)$ is often a spatiotemporal field for many engineering applications, and $\mathcal{G}^{\dagger}$ denotes the true dynamics, which may include differential and/or integral operators. If the detailed form of true dynamics is known, the BED can be used straightforwardly to estimate the unknown parameters $\bm{\theta}^{\dagger}$. In practice, the form of true dynamics is often unknown for many complex dynamical systems. Assuming that a physics-based model $\mathcal{G}(\mathbf{u};\bm{\theta}_{\mathcal{G}})$ exists to approximate the true dynamics $\mathcal{G}^{\dagger}$, we focus on the learning of model discrepancy $\mathcal{G}^{\dagger}-\mathcal{G}$, which is characterized by a neural network $\textrm{NN}(\mathbf{u};\bm{\theta}_{\textrm{NN}})$ and leads to the modeled system:
@@
Using a neural-network-based model to characterize the model discrepancy has recently been explored with different choices of neural networks such as neural operators [55] and diffusion models [56, 57], and with various types of research context such as ensemble data assimilation [58]. Although ill-posedness could generally be an issue for the joint learning of physics-based model and neural-network-based discrepancy, systematically preventing the neural network from ‘explaining away’ physically prescribed terms is not the main focus of this work, and the numerical examples studied in this work do not noticeably suffer from such an issue.
In the context of BED, the design $\mathbf{d}$ corresponds to the spatiotemporal coordinate $(\mathbf{z},t)$. The measurement $\mathbf{y}$ is the observation of variable $\mathbf{u}$ that represents the state of the physical system at the design $\mathbf{d}$, i.e., $\mathbf{y}=\mathbf{u}(\mathbf{d})+\bm{\eta}$, where $\bm{\eta}\sim\mathcal{N}(\bm{0},\bm{\Gamma})$ represents the measurement noise which is assumed to be Gaussian in this work when observing the state from the true system. More generally, the proposed framework can be extended to nonlinear observation systems, provided that the corresponding likelihood function can be explicitly defined based on the noise model.
@@ 2.1 Bayesian Experimental Design
where $\mathbf{y}\in\mathcal{Y}\subseteq\mathbb{R}^{d_{\mathbf{y}}}$ is data from the design $\mathbf{d}\in\mathcal{D}\subseteq\mathbb{R}^{d_{\mathbf{d}}}$, $\bm{\theta}\in\bm{\Theta}\subseteq\mathbb{R}^{\bm{\theta}}$ denotes the target parameters, and $U$ is the utility function taking the inputs of $\mathbf{y}$ and $\bm{\theta}$ and returning a real value. The optimal design $\mathbf{d}^{*}$ is obtained by maximizing the expected utility function $\mathbb{E}[U(\bm{\theta},\mathbf{y},\mathbf{d})]$ over the design space $\mathcal{D}$. The term $p(\bm{\theta},\mathbf{y}|\mathbf{d})$ is the joint conditional distribution of data and parameters. $p(\mathbf{y}|\bm{\theta},\mathbf{d})$ is the likelihood. One thing that should be noted is that $p(\bm{\theta}|\mathbf{d})$ is the prior distribution of parameters, which is often assumed to be independent of the design $\mathbf{d}$: $p(\bm{\theta}|\mathbf{d})=p(\bm{\theta})$.
The utility function $U(\bm{\theta},\mathbf{y},\mathbf{d})$ can be regarded as the reward obtained from the data $\mathbf{y}$ for the corresponding design $\mathbf{d}$. This work employs Kullback-Leibler divergence between the posterior and the prior distribution of parameters $\bm{\theta}$, which is a widely used utility function in BED [35] and has an information-theoretic interpretation[21, 28, 24]:
@@
In this work, the optimization problem is solved using a gradient-based approach, for which the gradient can be approximated via Infinitesimal Perturbation Analysis (IPA) [34]. In practice, it can also be computed using automatic differentiation. For the high-dimensional designs, comprehensive reviews of methods for solving design optimization problems can be found in [24, 36].
@@ 2.2 Model Discrepancy Correction
More specifically, the iterative approach at $i$-th stage starts with solving a BED problem for $\bm{\theta}_{\mathcal{G}}$ based on the current estimation of $\bm{\theta}_{\text{NN}}$ to obtain the optimal design $\mathbf{d}$ in Eq. (7). With the fixed design $\mathbf{d}$ and the neural network parameters $\bm{\theta}_{\text{NN}}$, we further obtain the maximum a posteriori (MAP) point $\bm{\theta}_{\mathcal{G}}^{*}$:
$\bm{\theta}_{\mathcal{G}}^{*}=\operatorname*{arg\,max}_{\bm{\theta}_{\mathcal{G}}}\{p(\mathbf{y}|\bm{\theta}_{\mathcal{G}},\mathbf{d},\bm{\theta}_{\text{NN}})p(\bm{\theta}_{\mathcal{G}}|\bm{\theta}_{\text{NN}})\}.$ (8)
where $p(\bm{\theta}_{\mathcal{G}}|\bm{\theta}_{\text{NN}})$ represents the prior distribution with $\mathbf{d}$ hiden since prior knowledge is independant with experimental design.
With the MAP estimation of $\bm{\theta}_{\mathcal{G}}^{*}$ further being fixed, the objective function $L(\bm{\theta}_{\text{NN}};\ \bm{\theta}_{\mathcal{G}}^{*},\mathbf{y},\mathbf{d})$ is defined by the likelihood function $p(\mathbf{y}|\bm{\theta}_{\mathcal{G}}^{*},\mathbf{d},\bm{\theta}_{\text{NN}})$ , and the updated estimation of $\bm{\theta}_{\text{NN}}$ can be obtained via:
@@
We omit the stage index $i$ in this section for simplicity in notation. In summary, there are three key steps at each stage of the hybrid framework illustrated in Fig. 1: (i) performing BED to obtain optimal design $\mathbf{d}$, (ii) obtaining the MAP estimation $\bm{\theta}_{\mathcal{G}}^{*}$, and (iii) updating the estimation of $\bm{\theta}_{\text{NN}}$. A sketch of proof for the existence of a globally optimal solution is presented in A. Here, the global optimum refers to the solution that the MAP estimate of $\bm{\theta}_{\mathcal{G}}$ aligns with the ground truth and the neural network perfectly recovers the discrepancy function. It is worth noting that gradient descent optimization methods would not always guarantee such a global optimal solution.
This iterative procedure exploits a distinction between the two types of parameters. We treat $\bm{\theta}_{\mathcal{G}}$ as a low-dimensional random vector inferred through Bayesian approaches. In contrast, the neural network parameters $\bm{\theta}_{\text{NN}}$ are optimized via gradient-based methods due to their high dimensionality, which makes a full Bayesian viewpoint computationally prohibitive. The above updating strategy resembles the Expectation-Maximization (EM) algorithm, in which one parameter set takes the expectation to facilitate the optimization of the other parameter set.
In practice, a single MAP estimation via Eq. (9) can sometimes lead to unsatisfactory performance. At intermediate stages of the sequential design, a limited amount of data can potentially lead to a posterior landscape with multiple, potentially misleading, local maxima. Any single one of these maxima could bias the algorithm and hinder subsequent decisions. To enhance the robustness of the proposed algorithm, we therefore seek to identify multiple local maxima from the posterior distribution of $\bm{\theta}_{\mathcal{G}}$:
$\{\bm{\theta}_{\mathcal{G}}^{(1)},\bm{\theta}_{\mathcal{G}}^{(2)},...,\bm{\theta}_{\mathcal{G}}^{(m)}\}=\operatorname{arg\,top}_{m}p(\mathbf{y}|\bm{\theta}_{\mathcal{G}},\mathbf{d},\bm{\theta}_{\text{NN}})p(\bm{\theta}_{\mathcal{G}}|\bm{\theta}_{\text{NN}}),$ (10)
where $\operatorname{arg\,top}_{m}$ denotes the process that identifies the set of parameters corresponding to the top $m$ largest values of a function. This can be approximated by solving the MAP problem in Eq. (8) $m$ times with randomly chosen initial points. The $\bm{\theta}_{\mathcal{G}}^{*}$ in Eq. (9) is then approximated by those samples (more related to posterior mean) :
@@
The primary purpose of this averaging is to produce a robust point estimate $\bm{\theta}_{\mathcal{G}}^{*}$ for the subsequent neural network update. While it is not guaranteed to be a local maximum of the posterior, numerical results suggest that the gradient it provides still benefits updating the neural network parameters. More specifically, the gradient of the objective function in Eq. (9) evaluated at the averaged point provides a descent direction that is balanced by the influence of all identified peaks. This mechanism prevents the network from overfitting to a single, potentially misleading, mode during the learning process. The limitations of this averaging strategy, particularly in cases where the ground truth posterior is genuinely multimodal, are further discussed in Section 4.
@@
It should be noted that $\mathrm{d}\mathbf{u}/\mathrm{d}\bm{\theta}_{\text{NN}}$ is often expensive to evaluate directly when $\mathbf{u}$ and $\bm{\theta}_{\text{NN}}$ are both high-dimensional. Therefore, efficient methods have been developed to obtain $\mathrm{d}\text{Obj}/\mathrm{d}\bm{\theta}_{\text{NN}}$ (i.e., $\partial L/\partial\bm{\theta}_{\text{NN}}$) such as automatic differentiation [63, 64] or adjoint methods [65, 66].
As automatic differentiation is essentially a discretized version of the adjoint method, we only briefly introduce the application of the adjoint method here. The brief introduction below is based on data at a given time step $t_{1}$, while derivations and extensions for handling time-integrated objective functions and data at multiple time steps can be found in [67, 66].
To avoid directly calculating $\mathrm{d}\mathbf{u}/\mathrm{d}\bm{\theta}_{\text{NN}}$, an adjoint state $\lambda(t)$ is defined such that $\lambda(t)^{\top}=\partial L/\partial\mathbf{u}(t)$. For the modeled system in Eq. (2), the adjoint state in time range $t_{0},t_{1}]$ can be obtained by backwardly solving:
$\left\{\begin{array}[]{l}\begin{aligned} \frac{\mathrm{d}\lambda(t)}{\mathrm{d}t}&=-\left(\frac{\partial(\mathcal{G}+\text{NN})}{\partial\mathbf{u}}\right)^{\top}\lambda(t)\\ \lambda^{\top}(t_{1})&=\frac{\partial L}{\partial\mathbf{u}(t_{1})}\end{aligned}\end{array}\right$., (13)
where $\partial\text{NN}/\partial\mathbf{u}$ can be obtained directly via the backpropagation of the neural network NN and $\partial\mathcal{G}/\partial\mathbf{u}$ is the Jacobian of the physics-based model $\mathcal{G}$. In practice, the Jacobian can be derived analytically or approximated by the coefficient matrix from the numerical solver, which often formulates a linear system. With the adjoint state obtained, the gradient $\mathrm{d}L/\mathrm{d}\bm{\theta}_{\text{NN}}$ can then be calculated as follows:
$\frac{\mathrm{d}L}{\mathrm{d}\bm{\theta}_{\text{NN}}}$ $=\int_{t_{0}}^{t_{1}}\lambda^{\top}(t)\frac{\partial\left(\mathcal{G}+\text{NN}\right)}{\partial\bm{\theta}_{\text{NN}}}dt$ (14)
$=\underbrace{\int_{t_{0}}^{t_{1}}\lambda^{\top}(t)\frac{\partial\mathcal{G}}{\partial\bm{\theta}_{\text{NN}}}dt}_{\mathclap{=0}}+\int_{t_{0}}^{t_{1}}\lambda^{\top}(t)\frac{\partial\text{NN}}{\partial\bm{\theta}_{\text{NN}}}dt,$
where the first term at the right-hand side has to be zero, as the physics-based model $\mathcal{G}$ does not depend on $\theta_{\text{NN}}$.